OpenStudy (anonymous):

Solve for x: 3|x − 3| + 2 = 14

OpenStudy (mathlegend):

So first I would change the equation to: 3(x-3)+2=14

OpenStudy (mathlegend):

@Guez can you solve that?

OpenStudy (mathlegend):

Step by Step.

OpenStudy (mathlegend):

@drastic please do not give answers..

OpenStudy (anonymous):

u can plug in and check that but i am sure that it is the right answer

OpenStudy (mathlegend):

@drasticc122 please read the code of conduct. He wants help and I am going to help him. Not give answers.

OpenStudy (anonymous):

sorry i am taking a math test right now but i just wanted to help him

OpenStudy (anonymous):

There will be 2 answers. leave the absolute value to the end. Do not get rid of it!! The main objective when you see these problems is to get the absolute value by itself on one side of the equation first. This is out first step. Same algebra principles apply

OpenStudy (anonymous):

3|x − 3| + 2 = 14 (subtract 2 from each side) 3|x − 3| = 12 (divide both sides by 3) |x − 3| = 4

OpenStudy (anonymous):

cheme is right

OpenStudy (anonymous):

|x − 3| = 4 Now what this means is the positive of x-3 = 4 and the negative of x-3 = 4. we take |x − 3| = 4 and write 2 equations, both giving us a solution for x. -(x-3) = 4 and +(x-3)=4 -x + 3 = 4 and x - 3 = 4 -x = 1 and x = 7 x = -1 and 7